OpenStudy (kaylala):

Find the Exact Value of: sin (-π/12) by using π/3 and π/4

OpenStudy (kaylala):

@mustafa2014

OpenStudy (callisto):

(-π/12) = π/4 - π/3

OpenStudy (callisto):

Also, sin(A-B)=sin A cos B - cos A sin B

OpenStudy (kaylala):

sin 45 cos 60 - cos 45 sin 50 ?

OpenStudy (kaylala):

sin 45 cos 60 - cos 45 sin 60 ?

OpenStudy (anonymous):

sin(-)=..... sin(a + b)=........

OpenStudy (callisto):

@kaylala Yes.

OpenStudy (kaylala):

((sqrt2)/4) - ((sqrt6)/4) @Callisto

OpenStudy (callisto):

Looks good :)

OpenStudy (kaylala):

okay thanks @Callisto